OpenStudy (anonymous):

If c is the measure of the hypotenuse of a right triangle, find each missing measure. If necessary, round to the nearest hundredth. a = 60, b = ?, c = 100 Select one: a. 80.00 b. 40 c. 160 d. 116.62

OpenStudy (haseeb96):

D is the correct answer

OpenStudy (anonymous):

haha no its not

OpenStudy (haseeb96):

yes it is you can check it by yourself apply phythagoras theorem (Hyp:)^2= (base)^2 + (perp:) ^2 c^2 = 60^2 + 100^2 c^2 =13600 taking square root on both sides c = 116.62

OpenStudy (haseeb96):

@AWillie98 did you get it?

OpenStudy (acxbox22):

hahaha thats wrong

OpenStudy (anonymous):

@acxbox22 thats what i told him

OpenStudy (acxbox22):

he didnt read the part where it says c=hypotenuse

OpenStudy (acxbox22):